USING uncertainity principle ,estimate the minimum energy of a particle in a simple harmonic potential U= 1/2 kx^2
1
Expert's answer
2020-01-06T10:16:49-0500

The uncertainty principle introduced first in 1927, by the German physicist Werner Heisenberg. Most often this principle is written as

(1) "\Delta x\cdot \Delta p_x \ge \frac{\hbar}{2}" , where "\Delta x" and "\Delta p_x" - standard deviation of position "x" and the standard deviation of momentum component "p_x" of particle under consideration, "\hbar =\frac{h}{2\pi}, h -"Plank's constant. According to the uncertainty principle, a simple harmonic oscillator can never be at rest. Its energy consists of the sum of the potential and kinetic energies of a particle with mass "m" moves in a simple harmonic potential "U(x)=\frac{k\cdot x^2}{2}" . That is

(2) "E_{sum}=\bar E_p+\bar E_k"

In the case of the lowest energy level, the standard square deviation coincides with the mean square of the coordinate "\bar {x^2}={\Delta x}^2" A similar conclusion can be made for the average square of the momentum "\bar {{p_x}^2}={\Delta p_x}^2" . Therefore (2) we can rewrite as (3)

(3) "E_0=\frac{k\cdot {\Delta x}^2}{2}+\frac{{\Delta p_x}^2}{2m}"

Assuming we have achieved exact equality in (1) and substituting "\Delta p_x =\frac{\hbar}{2\Delta x}" to (3) we get an expression for energy containing only "\Delta x".

(4) "E_0=\frac{k\cdot {\Delta x}^2}{2}+\frac{{\hbar}^2}{8m \cdot {\Delta x}^2}"

Differentiating (4) by "\Delta x" we find

"E_0^{'}=k\cdot \Delta x - \frac{{\hbar}^2}{4m \cdot {\Delta x}^3}"

The position of the minimum zero energy is

"E_0^{'}=0 \Rightarrow k\cdot \Delta x - \frac{{\hbar}^2}{4m \cdot {\Delta x}^3}=0 \Rightarrow \Delta x= ({\frac{{\hbar}^2}{4mk}})^{\frac{1}{4}} \Rightarrow {\Delta x}^2= \frac{{\hbar}}{\sqrt{4mk}}"


and the minimum energy value given by the formula (5) (substitute "{\Delta x}^2" to (4))

(5) "E_0=\frac{k\cdot {\frac{\hbar}{\sqrt{4mk}}}}{2}+\frac{{\hbar}^2}{8m \cdot {\frac{\hbar}{\sqrt{4mk}}}}=
\frac{ \hbar}{4}\cdot \sqrt{\frac{k}{m}}+\frac{ \hbar}{4}\cdot \sqrt{\frac{k}{m}}=\frac{\hbar}{2}\cdot \sqrt{\frac{k}{m}}"

It is known from the classical solution of the simple harmonic oscillator problem that "\sqrt{\frac{k}{m}}=\omega" the frequency of oscillation ("rad \cdot s^{-1}"). We can write

"E_0=\frac{\hbar \omega}{2}"